# Asymptotic Normality of Subgraph Counts in Sparse Inhomogeneous Random Graphs

### v1

@@ 2.4.3. Dense Graphon Model
Let $H$ be a subgraph with $|E(H)|\geq 1$ and t(H,W)>0.t(H,\,W)>0. Let $F\subseteq H$ be a subgraph of $H$ with $|V(F)|\geq 1.$ Then there exist copies $H_{s},\,H_{t}\in\mathcal{H}_{n}$ of $H$ in $K_{n}$ such that $H_{s}\cap H_{t}\simeq F$ and $t(H_{s}\cup H_{t},\,W)\geq(t(H,\,W))^{2}>0.$ Moreover, the number of such pairs $H_{s},\,H_{t}$ is $O_{H}(n^{2|V(H)|-|V(F)|}).$
Fix a subgraph $F\subseteq H$ such that $|V(F)|\geq 1.$ By Lemma 3.2, there are $O_{H}(n^{2|V(H)|-|V(F)|})$ pairs $H_{s},\,H_{t}\in\mathcal{H}_{n}$ such that $H_{s}\cap H_{t}\simeq F$ and $t(H_{s}\cup H_{t},\,W)\geq(t(H,\,W))^{2}>0.$ This means, when $|E(F)|\geq 1,$
$\mathrm{Cov}[I_{s},\,I_{t}]$ $=\mathbb{E}[I_{s}I_{t}]-(\rho_{n}^{|E(H)|}t(H,\,W))^{2}$
$=\rho_{n}^{2|E(H)|-|E(F)|}t(H_{s}\cup H_{t},\,W)-\rho_{n}^{2|E(H)|}(t(H,\,W))^{2}\gtrsim_{H,\,W}\rho_{n}^{2|E(H)|-|E(F)|},$
since $\rho_{n}\ll 1.$ Also, when $|E(F)|=0,$ by Lemma 3.2, we have $t(H_{s}\cup H_{t},\,W)\geq(t(H,\,W))^{2}.$ and, hence,
$\mathrm{Cov}[I_{s},\,I_{t}]$ $=\mathbb{E}[I_{s}I_{t}]-(\rho_{n}^{|E(H)|}t(H,\,W))^{2}$
$=\rho_{n}^{2|E(H)|}(t(H_{s}\cup H_{t},\,W)-(t(H,\,W))^{2})\gtrsim_{H,\,W}\rho_{n}^{2|E(H)|-|E(F)|}.$
Thus, (3.4) leads to
$\mathrm{Var}[X(H,G_{n})]$ $\gtrsim_{H,W}\sum_{F\subseteq H:|V(F)|\geq 1}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}$
$\gtrsim_{H,W}\max_{F\subseteq H:|V(F)|\geq 1}\frac{n^{2|V(H)|}\rho_{n}^{2|E(H)|}}{n^{|V(F)|}\rho_{n}^{|E(F)|}}.$ (3.7)
Combining (3.6) and (3.7), the result in (3.2) follows. $\Box$
@@
The proof of Proposition 2.15 will use Stein’s method based on dependency graphs. We begin by defining the notion of a conditional dependency graph, given a sigma-algebra.
@@
Let $\{X_{v}\}_{v\in V}$ be a family of random variables (defined on some common probability space) indexed by a finite set $V$ and $\mathcal{F}$ be a $\sigma$-algebra. Then a graph $\mathcal{G}$ with vertex set $V$ is said to be a conditional dependency graph for the collection $\{X_{v}\}$ given $\mathcal{F}$ if the following holds: For any two subsets of vertices $A,B\subseteq V$ such that there is no edge from in $\mathcal{G}$ from any vertex in $A$ to any vertex in $B,$ then the colelctions $\{X_{v}\}_{v\in A}$ and $\{X_{v}\}_{v\in B}$ are conditionally independent given $\mathcal{F}.$
@@
Proposition 4.2 ([34, Theorem 6.33]).
Suppose $\{X_{v}\}_{v\in V}$ be a family of random variables with conditional dependency graph $\mathcal{G}$ given $\mathcal{F}.$ Assume $\mathbb{E}(X_{v}|\mathcal{F})=0$ almost surely, for all $v\in V.$ Let
@@
Technically, [34, Theorem 6.33] establishes the above result in the unconditional setting. The conditional result stated above follows by repeating the proof verbatim (with expectations replaced by conditional expectations).
@@
$X_{s}=I_{s}-\mathbb{E}(I_{s}\mid\mathcal{F}(\bm{U}_{n}))=I_{s}-\rho_{n}^{|E(H)|}\prod_{(a,\,b)\in E(H_{s})}W(U_{a},\,U_{b}).$
Clearly, $\mathbb{E}[X_{s}\mid\mathcal{F}]=0$, for all $1\leq s\leq M$. Recalling the definition from $\Delta_{1}(H,\,G_{n})$ (1.5), note that
@@
Construct a conditional dependency graph $\mathcal{G}$ of the collection of random variables $\{X_{s}:H_{s}\in\mathcal{H}_{n}\}$ given $\sigma(\bm{U}_{n})$ on the vertex set $\{1,\,2,\,\ldots,\,M\}$ as follows: Connect the edge (s,t)(s,\,t), for $1\leq s\leq M$ in $\mathcal{G}_{n}$ if and only if $|E(H_{s})\cap E(H_{t})|\geq 1$.
@@
We have
$\sum_{s=1}^{M}\mathbb{E}[|X_{s}|\mid\mathcal{F}(\bm{U}_{n})]\lesssim_{H,\,W}\mathbb{E}[X(H,G_{n})],$ (4.2)
@@
[…]
almost surely. Further, with $M=|\mathcal{H}_{n}|=\binom{n}{|V(H)|}\frac{|V(H)|!}{|\mathrm{Aut}(H)|}\lesssim_{H,\,W}n^{|V(H)|}$,
$\sum_{s=1}^{M}\mathbb{E}[|X_{s}|\mid\mathcal{F}(\bm{U}_{n})]\lesssim_{H,\,W}n^{|V(H)|}\rho_{n}^{|E(H)|}\asymp_{H,\,W}\mathbb{E}[X(H,\,G_{n})]$ (4.4)
@@
$\sum_{w\in\bar{N}_{\mathcal{G}}(s,\,t)}\mathbb{E}[|X_{w}|\mid X_{s},\,X_{t},\,\mathcal{F}(\bm{U}_{n})]\lesssim_{H,W}\frac{\mathbb{E}[X(H,G_{n})]}{\Phi_{H}},$ (4.5)
@@
$\Phi_{H}:=\min\{\mathbb{E}[X(F,\,G_{n})]:F\subseteq H,\,|V(F)|\geq 1\}\asymp\min_{F\subseteq H,\,|V(F)|\geq 1}n^{|V(F)|}\rho_{n}^{|E(F)|}.$
@@
since $\rho_{n}\ll 1$. If we fix a subgraph $F\subseteq H,$ then the number of choices of $w\in\{1,\,2,\,\ldots,\,M\}$ such that $H_{w,\,\{s,\,t\}}$ is isomorphic to $F$ is $O_{H}(n^{|V(H)|-|V(F)|})$. Furthermore, note that $w\in\bar{N}_{\mathcal{G}}(s,\,t)$ if and only if $|E(H_{w,\,\{s,\,t\}})|>0.$ Hence, due to (4), (4.5) can be bounded as:
@@
$\lesssim_{H,W}\max_{F\subseteq H:|V(F)|\geq 1}n^{|V(H)|-|V(F)|}\rho_{n}^{|E(H)|-|E(F)|}$
$\asymp_{H,\,W}\frac{\mathbb{E}[X(H,\,G_{n})]}{\Phi_{H}},$
@@
Suppose $\frac{\mathrm{Var}[\Delta_{1}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\to\kappa$ for some constant $\kappa>0$. Then
$(\sigma(\mathcal{F}(\bm{U}_{n})))^{2}\gtrsim_{H,W}\frac{(\mathbb{E}[X(H,G_{n})])^{2}}{\Phi_{H}},$ (4.8)
@@
Observe that $Z_{n}(F)$ and $A_{n}(F)$ are $U$-statistics with kernels of order 2|V(H)||V(F)|.2|V(H)|-|V(F)|. Hence, from the strong law for $U$-statistics (see [16, Theorem 6.3.1]) it follows that $\frac{Z_{n}(F)}{\mathbb{E}[Z_{n}(F)]}\to 1$ and $\frac{A_{n}(F)}{\mathbb{E}[A_{n}(F)]}\to 1$ almost surely. Thus,
@@
since there are finitely many terms in the sums in both numerator and denominator and for every fixed $F,$ $\frac{Z_{n}(F)}{\mathbb{E}[Z_{n}(F)]}\to 1$ and $\frac{A_{n}(F)}{\mathbb{E}[A_{n}(F)]}\to 1$, almost surely. Since we assumed that $\frac{\mathrm{Var}[\Delta_{1}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\to\kappa$ for some constant $\kappa>0,$ using (3.2) gives,
$\mathrm{Var}[\Delta_{1}(H,\,G_{n})]\asymp_{H,W}\mathrm{Var}[\Delta(H,G_{n})]$ $\asymp_{H,W}\max_{F\subseteq H:|V(F)|\geq 1}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}$
$\asymp\frac{(\mathbb{E}[X(H,G_{n})])^{2}}{\Phi_{H}}$ (4.12)
due to Lemma 3.1. Combining (4.11) and (4.12), the result in (4.8) follows. ∎
@@
$T(H,\,G_{n})=\frac{1}{\binom{n}{|V(H)|}}\sum_{1\leq i_{1}<\ldots<i_{|V(H)|}\leq n}h(U_{i_{1}},\,U_{i_{2}},\,\ldots,\,U_{i_{|V(H)|}}),$ (5.1)
@@
$\Delta_{2}(H,\,G_{n})$ $=\binom{n}{|V(H)|}\rho_{n}^{|E(H)|}T(H,\,G_{n})$ (5.2)
Note that $T(H,\,G_{n})$ is a centered $U$-statistics of order $|V(H)|$ with kernel $h.$ In the next lemma we show that the variance of $T(H,G_{n})$ scales as $\frac{1}{n}$ when the graphon $W$ is $H$-regular. The proof is deferred to Appendix A.3.
@@
Theorem 2.6 assumes that $W$ is $H$-regular. Now recalling (5.2), and using Lemma 5.1,
@@
This means, recalling (3.2),
$\frac{\mathrm{Var}[\Delta_{2}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\lesssim_{H,W}\min_{F\subseteq H:|V(F)|\geq 1}n^{|V(F)|-2}\rho_{n}^{|E(F)|}\rightarrow 0.$ (5.3)
This proves the first assertion in (2.5). Next, using (2.3), $\frac{\mathrm{Var}[\Delta_{1}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}=1-\frac{\mathrm{Var}[\Delta_{2}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\to 1$. This means, from (4.11), $\frac{\mathrm{Var}[\Delta_{1}(H,\,G_{n})|\mathcal{F}(\bm{U}_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\rightarrow 1$. Hence, applying Proposition 2.15 and Slutsky’s theorem it follows that
@@
In this section, we prove Theorem 2.11, which considers the case where $W$ is $H$-regular. We begin by showing that $\Delta_{2}(H,\,G_{n})$ (respectively, $\Delta_{1}(H,\,G_{n})$) has negligible contribution to the total variance when $n\rho_{n}^{m_{1}(H)}\ll 1$ (respectively, $n\rho_{n}^{m_{1}(H)}\gg 1$).

### v2

@@ 2.4.3. Dense Graphon Model
Let $H$ be a subgraph with $|E(H)|\geq 1$ and $t(H,W)>0.$ Let $F\subseteq H$ be a subgraph of $H$ with $|V(F)|\geq 1.$ Then there exist copies $H_{s},H_{t}\in\mathcal{H}_{n}$ of $H$ in $K_{n}$ such that $H_{s}\cap H_{t}\simeq F$ and $t(H_{s}\cup H_{t},W)\geq t(H,W)^{2}>0.$ Moreover, the number of such pairs $H_{s},H_{t}$ is $\asymp_{H}(n^{2|V(H)|-|V(F)|}).$
We begin by decomposing (3.6) according to the following 3 parts:
$\mathrm{Var}[X_{n}(H,G_{n})]$
$=$ $\,\sum_{F\subseteq H:|V(F)|=1}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}[\rho_{n}^{2|E(H)|-|E(F)|}t(H_{s}\cup H_{t},W)-(\rho_{n}^{|E(H)|}t(H,W))^{2}]$
$\,+\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|=0\end{subarray}}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}[\rho_{n}^{2|E(H)|-|E(F)|}t(H_{s}\cup H_{t},W)-(\rho_{n}^{|E(H)|}t(H,W))^{2}]$
$\,+\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|\geq 1\end{subarray}}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}[\rho_{n}^{2|E(H)|-|E(F)|}t(H_{s}\cup H_{t},W)-(\rho_{n}^{|E(H)|}t(H,W))^{2}]$
$=:$ $\,T_{1}+T_{2}+T_{3}.$ (3.7)
We begin by considering $T_{1}$. Since in this case $|E(F)|=0,$ we may factor out $\rho_{n}^{2|E(H)|}$ to obtain
$T_{1}=\rho_{n}^{2|E(H)|}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq K_{1}\end{subarray}}[t(H_{s}\cup H_{t},W)-t(H,W)^{2}],$ (3.8)
where $K_{1}$ denotes a single isolated vertex. Now, to identify the structure of $H_{s}\cup H_{t},$ we need the following definition:
Definition 3.3.
Suppose $H=(V(H),E(H))$ with vertices labeled $V(H)=\{1,2,\ldots,|V(H)|\}$. Then for $1\leq a,b\leq|V(H)|$, the $(a,b)$-vertex join of two copies of $H,$ denoted by $H\bigoplus_{a,b}H$, is the graph obtained by identifying the $a$-th vertex of the first copy $H$ with the $b$-th vertex of the second copy $H$ (see Figure 3).
Figure 3. The $(a,b)$-vertex join of the 2 copies of $H$.
Note that if $H_{s},H_{t}\in\mathcal{H}_{n}$ such that $H_{s}\cap H_{t}\simeq K_{1},$ then $H_{s}\cup H_{t}$ is isomorphic to $H\bigoplus_{a,b}H$ for some choice of $a,b\in V(H).$ Also, observe that for each fixed pair $(a,b)$, the number of ordered pairs $(H_{s},H_{t})$ such that $H_{s}\cap H_{t}\simeq K_{1}$ is $\asymp_{H}n^{2|V(H)|-1}$. Hence,
$T_{1}$ $=\rho_{n}^{2|E(H)|}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq K_{1}\end{subarray}}[t(H_{s}\cup H_{t},W)-t(H,W)^{2}]$
$\asymp_{H,W}n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}\left[\sum_{1\leq a,b\leq|V(H)|}t\left(H\bigoplus_{a,b}H,W\right)-|V(H)|^{2}t(H,W)^{2}\right].$
By Bhattacharya et al. (2023, Lemma 5.4) the bracketed term in the above expression is zero if and only if $W$ is $H$-regular. Thus,
$T_{1}\asymp_{H,W}\begin{cases}n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}&\text{if $W$ is not $H$-regular,}\\ 0&\text{if $W$ is $H$-regular.}\end{cases}$ (3.9)
Next, we consider $T_{2}$. Since $|E(F)|=0,$ we again factor out $\rho_{n}^{2|E(H)|}$ to get,
$T_{2}=\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|=0\end{subarray}}\rho_{n}^{2|E(H)|}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}[t(H_{s}\cup H_{t},W)-t(H,W)^{2}].$
For a fixed subgraph $F\subseteq H,$ the number of ordered pairs $(H_{s},H_{t})$ with $H_{s}\cap H_{t}\simeq F$ is $\asymp_{H}n^{2|V(H)|-|V(F)|}.$ Hence,
$|T_{2}|\lesssim_{H,W}\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|=0\end{subarray}}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|}\lesssim_{H,W}n^{2|V(H)|-2}\rho_{n}^{2|E(H)|}.$ (3.10)
Finally, we consider $T_{3}$:
$T_{3}=\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|\geq 1\end{subarray}}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}[\rho_{n}^{2|E(H)|-|E(F)|}t(H_{s}\cup H_{t},W)-(\rho_{n}^{|E(H)|}t(H,W))^{2}].$ (3.11)
Recall that for each fixed subgraph $F\subseteq H,$ the number of ordered pairs $(H_{s},H_{t})$ with $H_{s}\cap H_{t}\simeq F$ is $\asymp_{H}n^{2|V(H)|-|V(F)|}.$ Therefore, if dropping the second term inside the bracket in the above expression and upper bound each of the homomorphism densities $t(H_{s}\cup H_{t},W)$ by 1 gives,
$T_{3}\lesssim_{H,W}\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|\geq 1\end{subarray}}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}.$ (3.12)
For the matching lower bound, fix $F\subseteq H$ with $|E(F)|\geq 1.$ By Lemma 3.2, for each subgraph $F\subseteq H,$ we can find $\asymp_{H}n^{2|V(H)|-|V(F)|}$ pairs $(H_{s},H_{t})$ with $H_{s}\cap H_{t}\simeq F$ and $t(H_{s}\cup H_{t},W)>0.$ Restricting the inner sum in (3.11) to these pairs, the first term inside the bracket in (3.11) is strictly positive. Moreover, since $\rho_{n}\ll 1$ and $|E(F)|\geq 1,$
$(\rho_{n}^{|E(H)|}t(H,W))^{2}\ll\rho_{n}^{2|E(H)|-|E(F)|}.$
Consequently,
$T_{3}\gtrsim_{H,W}\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|\geq 1\end{subarray}}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}.$ (3.13)
Combining (3.12) and (3.13), we conclude
$T_{3}$ $\asymp_{H,W}\sum_{\begin{subarray}{c}F\subseteq H:|V(F)|\geq 2,\\ |E(F)|\geq 1\end{subarray}}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}$
$\asymp_{H,W}\max_{F\subseteq H:|E(F)|\geq 1}\frac{n^{2|V(H)|}\rho_{n}^{2|E(H)|}}{n^{|V(F)|}\rho_{n}^{|E(F)|}}.$ (3.14)
Note that, recalling (3.10),
$\frac{|T_{2}|}{T_{3}}\asymp_{H,W}\min_{F\subseteq H:|E(F)|\geq 1}n^{|V(F)|-2}\rho_{n}^{|E(F)|}\lesssim_{H}\rho_{n}\ll 1.$
Hence, $|T_{2}|\ll T_{3}$ and combining (3.9) and (3.14) with (3.7) gives the following:
  • If $W$ is not $H$-regular, then
    $\mathrm{Var}[X_{n}(H,G_{n})]$ $\asymp_{H,W}n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}+\max_{F\subseteq H:|E(F)|\geq 1}\frac{n^{2|V(H)|}\rho_{n}^{2|E(H)|}}{n^{|V(F)|}\rho_{n}^{|E(F)|}}$
    $\asymp_{H,W}\max_{F\subseteq H:|V(F)|\geq 1}\frac{n^{2|V(H)|}\rho_{n}^{2|E(H)|}}{n^{|V(F)|}\rho_{n}^{2|E(F)|}}.$
  • If $W$ is $H$-regular,
    $\mathrm{Var}[X_{n}(H,G_{n})]$ $\asymp_{H,W}\max_{F\subseteq H:|E(F)|\geq 1}\frac{n^{2|V(H)|}\rho_{n}^{2|E(H)|}}{n^{|V(F)|}\rho_{n}^{|E(F)|}}.$
This completes the proof of (3.4). $\Box$
@@
The proof of Proposition 2.15 will use Stein’s method based on dependency graphs. We begin by defining the notion of a conditional dependency graph.
@@
Let $\{X_{v}\}_{v\in V}$ be a family of random variables (defined on some common probability space) indexed by a finite set $V$ and $\mathcal{F}$ be a $\sigma$-algebra. Then a graph $\mathcal{G}$ with vertex set $V$ is said to be a conditional dependency graph for the collection $\{X_{v}\}$ given $\mathcal{F}$ if the following holds: For any two subsets of vertices $A,B\subseteq V$ such that there is no edge in $\mathcal{G}$ from any vertex in $A$ to any vertex in $B,$ then the colelctions $\{X_{v}\}_{v\in A}$ and $\{X_{v}\}_{v\in B}$ are conditionally independent given $\mathcal{F}.$
@@
Proposition 4.2 ((Janson et al., 2011, Theorem 6.33)).
Suppose $\{X_{v}\}_{v\in V}$ be a family of random variables with conditional dependency graph $\mathcal{G}$ given $\mathcal{F}.$ Assume $\mathbb{E}[X_{v}|\mathcal{F}]=0$ almost surely, for all $v\in V.$ Let
@@
Technically, Janson et al. (2011, Theorem 6.33) establishes the above result in the unconditional setting. The conditional result stated above follows by repeating the proof verbatim (with expectations replaced by conditional expectations).
@@
$X_{s}:=I_{s}-\mathbb{E}[I_{s}\mid\mathcal{F}(\bm{U}_{n})]=I_{s}-\rho_{n}^{|E(H)|}\prod_{(a,b)\in E(H_{s})}W(U_{a},U_{b}).$
Clearly, $\mathbb{E}[X_{s}\mid\mathcal{F}]=0$, for all $1\leq s\leq M$. Recalling the definition of $\Delta_{1}(H,G_{n})$ from (1.5), note that
@@
Construct a conditional dependency graph $\mathcal{G}$ of the collection of random variables $\{X_{s}:H_{s}\in\mathcal{H}_{n}\}$ given $\sigma(\bm{U}_{n})$ on the vertex set $\{1,2,\ldots,M\}$ as follows: Connect the edge $(s,t)$, for $1\leq s\neq t\leq M$ in $\mathcal{G}_{n}$ if and only if $|E(H_{s})\cap E(H_{t})|\geq 1$.
@@
For $X_{s}$ as defined above,
$\sum_{s=1}^{M}\mathbb{E}[|X_{s}|\mid\mathcal{F}(\bm{U}_{n})]\lesssim_{H,W}\mathbb{E}[X_{n}(H,G_{n})],$ (4.2)
@@
$\lesssim_{H,W}\rho_{n}^{|E(H)|},$ (4.3)
almost surely. Since $M=|\mathcal{H}_{n}|=\binom{n}{|V(H)|}\frac{|V(H)|!}{|\mathrm{Aut}(H)|}\lesssim_{H,W}n^{|V(H)|}$,
$\sum_{s=1}^{M}\mathbb{E}[|X_{s}|\mid\mathcal{F}(\bm{U}_{n})]\lesssim_{H,W}n^{|V(H)|}\rho_{n}^{|E(H)|}\asymp_{H,W}\mathbb{E}[X_{n}(H,G_{n})],$ (4.4)
@@
$\sum_{w\in\bar{N}_{\mathcal{G}}(s,t)}\mathbb{E}[|X_{w}|\mid X_{s},X_{t},\mathcal{F}(\bm{U}_{n})]\lesssim_{H,W}\frac{\mathbb{E}[X_{n}(H,G_{n})]}{\Phi_{H}},$ (4.5)
@@
$\Phi_{H}:=\min\{\mathbb{E}[X_{n}(F,G_{n})]:F\subseteq H,|E(F)|\geq 1\}\asymp\min_{F\subseteq H,\,|E(F)|\geq 1}n^{|V(F)|}\rho_{n}^{|E(F)|}.$
@@
since $\rho_{n}\ll 1$. If we fix a subgraph $F\subseteq H,$ then the number of choices of $w\in\{1,2,\ldots,M\}$ such that $H_{w,\{s,t\}}$ is isomorphic to $F$ is $O_{H}(n^{|V(H)|-|V(F)|})$. Furthermore, note that $w\in\bar{N}_{\mathcal{G}}(s,t)$ if and only if $|E(H_{w,\{s,t\}})|>0.$ Hence, using (4), the LHS of (4.5) can be bounded as:
@@
$\asymp_{H,W}\max_{F\subseteq H:|E(F)|\geq 1}n^{|V(H)|-|V(F)|}\rho_{n}^{|E(H)|-|E(F)|}$
$\asymp_{H,W}\frac{\mathbb{E}[X_{n}(H,G_{n})]}{\Phi_{H}},$
@@
For $\sigma(\mathcal{F}(\bm{U}_{n}))$ as defined before, we have
$(\sigma(\mathcal{F}(\bm{U}_{n})))^{2}\gtrsim_{H,W}\frac{(\mathbb{E}[X_{n}(H,G_{n})])^{2}}{\Phi_{H}},$ (4.8)
@@
Observe that $Z_{n}(F)$ and $A_{n}(F)$ are scaled $U$-statistics with kernels of order 2|V(H)||V(F)|.2|V(H)|-|V(F)|. Hence, from the strong law for $U$-statistics (see Bose et al. (2024, Theorem 6.3.1)) it follows that $\frac{Z_{n}(F)}{\mathbb{E}[Z_{n}(F)]}\to 1$ and $\frac{A_{n}(F)}{\mathbb{E}[A_{n}(F)]}\to 1$ almost surely. Thus,
@@
since there are finitely many terms in the sums in both numerator and denominator and for every fixed $F,$ $\frac{Z_{n}(F)}{\mathbb{E}[Z_{n}(F)]}\to 1$ and $\frac{A_{n}(F)}{\mathbb{E}[A_{n}(F)]}\to 1$, almost surely. Now,
$\mathrm{Var}[\Delta_{1}(H,G_{n})]=\sum_{F\subseteq H:|E(F)|\geq 1}\left\{\rho_{n}^{2|E(H)|-|E(F)|}\mathbb{E}[Z_{n}(F)]-\rho_{n}^{2|E(H)|}\mathbb{E}[A_{n}(F)]\right\}.$ (4.12)
Fix $F\subseteq H$ with $|E(F)|\geq 1.$ Recall that by Lemma 3.2, for each subgraph $F\subseteq H,$ we can find $\asymp_{H}(n^{2|V(H)|-|V(F)|})$ pairs $(H_{s},H_{t})$ with $H_{s}\cap H_{t}\simeq F$ and $t(H_{s}\cup H_{t},W)>0.$ Thus, by definition of $A_{n}(F)$ and $Z_{n}(F),$ we have
$\mathbb{E}[A_{n}(F)]\lesssim_{H,W}\mathbb{E}[Z_{n}(F)]\asymp_{H,W}n^{2|V(H)|-|V(F)|}.$
Moreover, since $\rho_{n}\ll 1$ and $|E(F)|\geq 1,$
$\rho_{n}^{2|E(H)|}\mathbb{E}[A_{n}(F)]\ll\rho_{n}^{2|E(H)|-|E(F)|}\mathbb{E}[Z_{n}(F)].$
Thus, the second term in the within the brackets in (4.12) is of smaller order than the first term. Consequently,
$\mathrm{Var}[\Delta_{1}(H,G_{n})]$ $\asymp_{H,W}\sum_{F\subseteq H:|E(F)|\geq 1}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}$
$\asymp_{H,W}\max_{F\subseteq H:|E(F)|\geq 1}\frac{n^{2|V(H)|}\rho_{n}^{2|E(H)|}}{n^{2|V(F)|}\rho_{n}^{2|E(F)|}}\asymp_{H,W}\frac{(\mathbb{E}[X(H,G_{n})])^{2}}{\Phi_{H}}.$ (4.13)
Combining (4.11) and (4.13), the result in (4.8) follows. ∎
@@
$T(H,G_{n}):=\frac{1}{\binom{n}{|V(H)|}}\sum_{1\leq i_{1}<\ldots<i_{|V(H)|}\leq n}h(U_{i_{1}},U_{i_{2}},\ldots,U_{i_{|V(H)|}}),$ (5.1)
@@
$\Delta_{2}(H,G_{n})$ $=\binom{n}{|V(H)|}\rho_{n}^{|E(H)|}T(H,G_{n}).$ (5.2)
Note that $T(H,G_{n})$ is a centered $U$-statistics of order $|V(H)|$ with kernel $h.$ In the next lemma we show that the variance of $T(H,G_{n})$ scales as $\frac{1}{n^{2}}$ when the graphon $W$ is $H$-regular. The proof is deferred to Appendix A.3.
@@
Theorem 2.6 assumes that $W$ is $H$-regular. Now, recalling (5.2), and using Lemma 5.1,
@@
This means, recalling (3.4) in the $H$-regular case,
$\frac{\mathrm{Var}[\Delta_{2}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\lesssim_{H,W}\min_{F\subseteq H:|E(F)|\geq 1}n^{|V(F)|-2}\rho_{n}^{|E(F)|}\lesssim\rho_{n}\rightarrow 0.$ (5.3)
This proves the first assertion in (2.5). Next, using (2.3),
$\frac{\mathrm{Var}[\Delta_{1}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}=1-\frac{\mathrm{Var}[\Delta_{2}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\to 1.$
This means, from (4.11), $\frac{\mathrm{Var}[\Delta_{1}(H,G_{n})|\mathcal{F}(\bm{U}_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\rightarrow 1$. Hence, applying Proposition 2.15 and Slutsky’s theorem it follows that
@@
In this section, we prove Theorem 2.11, which considers the case where $W$ is not $H$-regular. We begin by showing that $\Delta_{2}(H,G_{n})$ (respectively, $\Delta_{1}(H,G_{n})$) has negligible contribution to the total variance when $n\rho_{n}^{m_{1}(H)}\ll 1$ (respectively, $n\rho_{n}^{m_{1}(H)}\gg 1$).